# GENIE: Watermarking Graph Neural Networks for Link Prediction

### v1

@@ 5.4 Non-ownership piracy
$\mathcal{A}$ may attempt to follow GENIE to insert his pirated watermark into a model stolen from $\mathcal{O}$ (i.e., $\mathcal{M}_{wm}$). However, owing to the robustness of GENIE (as witnessed in § 5.3), $\mathcal{A}$ will not be able to remove $\mathcal{O}$’s original watermark from $\mathcal{M}_{wm}$. Consequently, $\mathcal{A}$ cannot derive $\mathcal{M}_{adv}$ from $\mathcal{M}_{wm}$, which solely contains his pirated watermark and not $\mathcal{O}$’s original watermark.
GENIE injects watermark into an untrained model while $\mathcal{A}$ has access to only $\mathcal{M}_{wm}$. In a real-world setting, $\mathcal{A}$ can still generate his own pirated trigger set (using the method explained in § 4.1) and train stolen $\mathcal{M}_{wm}$ on the pirated trigger set to obtain $\mathcal{M}_{adv}$ (generally called a pirated model). Given that training on just the pirated trigger set might lead to decrease in $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$, $\mathcal{A}$ would want to identify an optimal number of epochs for training with the pirated trigger set such that $\mathcal{M}_{adv}$ has high AUC on both $\mathcal{D}_{test}$ and the pirated trigger set. Figure 5 shows the variations in $\mathcal{M}_{adv}$’s performance on $\mathcal{D}_{wm}$, $\mathcal{D}_{test}$, and pirated trigger set during pirated watermark embedding process across different numbers of epochs for GCN over NS dataset; cf. Appendix D.3 for other datasets.
Figure 5: A representative example of $\mathcal{M}_{adv}$’s performance trajectory on $\mathcal{D}_{wm}$, $\mathcal{D}_{test}$, and pirated trigger set during embedding of pirated watermark across training epochs.
We see that $\mathcal{M}_{adv}$ performs well on $\mathcal{D}_{wm}$, $\mathcal{D}_{test}$, as well as on pirated trigger sets around $20^{th}$ epoch. If $\mathcal{O}$ challenges $\mathcal{A}$ to present his model at this point, $\mathcal{M}_{adv}$ will contain $\mathcal{A}$’s pirated watermark as well as $\mathcal{O}$’s watermark. However, $\mathcal{O}$ can present $\mathcal{M}_{wm}$ containing only her watermark. Thus, identifying the true owner will be easy in such a dispute. We further observe that around $250^{th}$ epoch, $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{adv}}$ drops below the watermark threshold (cf. TABLE III), but $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{adv}}$ falls below the tolerable utility loss (i.e., up to 10%; following the definition of failure in §5.1) $\mathcal{A}$ is willing to tolerate. Even if $\mathcal{A}$ chooses to train for even higher epochs while embedding the pirated watermark, $\mathcal{M}_{adv}$ continues to lose its utility; rendering the $\mathcal{M}_{adv}$ useless. Hence, we take the liberty to claim that GENIE is robust against piracy attacks (i.e., $\mathcal{A}$ cannot fraudulently claim ownership or fabricate watermark over a pirated model).
@@ 5.5 Efficiency
The computational efficiency of a watermarking scheme is a critical aspect since it directly affects its cost and practical adoption. We rigorously evaluate the efficiency of GENIE by comparing the time required for the standard model training with the time needed for watermark embedding (which includes the time dedicated to generate $\mathcal{D}_{wm}$). In practice, the overall computational overhead mainly depends on the percentage of samples we choose to create $\mathcal{D}_{wm}$. If the $\mathcal{D}_{wm}$ is too big, embedding watermark information might take longer and even lead to a considerable loss in utility. On the other side, if $\mathcal{D}_{wm}$ is too small, it would be easier for $\mathcal{A}$ to remove the watermark (for instance, via pruning, fine-tuning).
We empirically observed that using percentage of samples up to 40% is optimal for SEAL, GCN, and GraphSAGE. It is the same percentage of samples we used in all our evaluations. TABLE XV shows the time needed for the standard model training against the time required for GENIE’s watermark embedding. Our results indicate that GENIE has reasonable computational overhead in exchange of the invaluable IP protection it confers to a model. SEAL over Yeast dataset consumed the maximum additional training time, which is only 18.3% more than the standard training time. Therefore, we argue that GENIE is efficient.
TABLE XV: Standard vs. watermark training in seconds.
Dataset SEAL GCN GraphSAGE
$T_{std}$ $T_{wm}$ $T_{std}$ $T_{wm}$ $T_{std}$ $T_{wm}$
C.ele 197.073 294.520 8.723 11.969 9.740 13.650
USAir 197.822 261.690 8.464 11.940 9.357 13.694
NS 112.798 138.869 8.933 12.365 9.909 14.224
Yeast 576.529 682.085 13.146 15.704 14.008 17.634
Power 270.702 345.792 11.229 14.283 12.130 16.612
arXiv 2009.926 2072.909 64.396 70.534 67.059 75.979
PPI 1048.213 1108.650 23.485 24.500 24.648 26.136
6 Discussion, limitation, future work
Our thorough evaluations show that GENIE satisfies all the watermarking requirements (cf. §3), i.e., functionality preservation (cf. §5.2), Un-removability (cf. §5.3), non-ownership piracy (cf. §5.4), efficiency (cf. §5.5), non-trivial ownership (cf. §4.3), capacity requirement does not apply to it since GENIE is a zero-bit watermarking scheme, and we evaluated GENIE on 3 GNN model architectures and 7 real-world datasets for the generality requirement. Given the salient features of GENIE, especially its efficiency in accommodating large datasets and performance on diverse datasets of varying scales and sparsities, we can claim that GENIE is suitable for real-world applications.
Uniqueness: As discussed in §2.4, the two state-of-the-art backdoor attacks on LP [42, 43] are not practical for watermarking LP task in GNNs. To this end, we implement GENIE using a backdoor that employs watermark feature vector (as described in §4.1.1 and §4.1.2) for watermarking GNNs without Data Poisoning (i.e., merging $\mathcal{D}_{wm}$ to $\mathcal{D}_{train}$ at each epoch) [37]. As opposed to the conventional data poisoning-based backdoor attacks [16], where loss function remains untouched, we propose two different loss functions (viz., $\mathcal{L}_{train}$, $\mathcal{L}_{wm}$) and a unique method of optimizing them individually. In particular, we use the gradient update mechanism described in Eq. 8 for the loss function optimization [^8] When using the conventional data poisoning technique to inject $\mathcal{D}_{wm}$, we observe $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ only up to $25\%$ for arXiv and PPI datasets. However, when using $\mathcal{L}=\mathcal{L}_{train}+\mathcal{L}_{wm}$, we observe a drop from $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ to $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ of upto 7%. Finally, by updating our gradients via Eq. 8, we observe a drop from $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{clean}}$ to $AUC_{\mathcal{D}_{test}}^{\mathcal{M}_{wm}}$ only upto 2% while maintaining $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ close to $100\%$ in most cases.. Our unique improvements cater to the superior performance of our proposed watermark injection method.
Stealthiness: A crucial requirement for a watermarking scheme is its evasiveness to detection. §5.2 demonstrates that $\mathcal{M}_{clean}$ and $\mathcal{M}_{wm}$ are indistinguishable in terms of their behavior on $\mathcal{D}_{test}$, proving the stealthiness of GENIE. The behavior of these models differs only on $\mathcal{D}_{wm}$. It is worth highlighting that only $\mathcal{O}$ has access to $\mathcal{D}_{wm}$ while it is computationally infeasible for $\mathcal{A}$ to guess the trigger set.
Limitation: Despite extensive empirical evaluations of GENIE in our study, GENIE does not currently have the theoretical guarantees required to ensure that the watermark will not be removed by any new attack in the future. Nevertheless, we present a competitive ownership demonstration scheme that statistically guarantees FPR and FNR [^9] This is equivalent to the probability of a daily event happening once every 2,739 years. of GENIE to be less than $10^{-6}$.
Future work: We will evaluate GENIE with more GNN architectures to further ensure its generalizability. In this paper, we consider LP in a transductive setting, which is considered as the default setting for LP task. In the future, we will extend GENIE to an inductive setting. Finally, all the graphs considered in this work are static. In the future, we will explore the applicability of GENIE to dynamic graphs.
@@ 7 Conclusion
Despite the tremendous success of GNNs in learning graph-structured data, protecting trained GNN models from model-stealing attacks is a critical issue. Existing GNN watermarking schemes focus either on node or graph classification tasks. In this paper, we propose GENIE, a backdoor-based watermarking scheme for GNNs tailored to LP task. We design GENIE for node representation-based and subgraph-based methods of LP. Our extensive evaluations show that GENIE is functionality-preserving. At the same time, GENIE is robust against state-of-the-art backdoor defenses and model extraction attacks. The statistical guarantees given by the ownership demonstration process confirm a close to zero probability of misclassification in GENIE. We hope our work sets new research directions and benchmarks in the domain of model watermarking.
Appendix A Further details on our experiment setup
We use a learning rate of 0.0001 for SEAL and 0.001 for GCN and GraphSAGE. Both $\mathcal{L}_{train}$ and $\mathcal{L}_{wm}$ use the same loss function (i.e., negative log likelihood) and optimizer (i.e., Adam) with the same learning rate. TABLE A.1 lists our watermarking rate for each dataset and respective model.
TABLE A.1: Watermarking rate (%) used in our experiments.
@@ Appendix A Further details on our experiment setup
SEAL 30 30 35 20 40 - -
Appendix B Statistical guarantee of non-trivial ownership in GENIE

### v3

@@ 5.5 Non-ownership piracy
Ownership piracy attacks allow $\mathcal{A}$ to follow Genie to insert her pirated watermark into a model stolen from $\mathcal{O}$. This creates an ambiguity for $\mathcal{J}$ to determine which party has originally watermarked the given model. However, for such an attack to succeed, $\mathcal{A}$ must present $\mathcal{M}_{adv}$ containing only her pirated watermark, free of the $\mathcal{O}$’s original watermark. It is because if $\mathcal{M}_{adv}$ contains both $\mathcal{O}$’s and $\mathcal{A}$’s watermark, $\mathcal{J}$ can easily resolve the conflict by comparing the timestamps of both watermarks (cf. §4.4). However, owing to the robustness of Genie (as witnessed in § 5.4), $\mathcal{A}$ will not be able to remove $\mathcal{O}$’s original watermark from $\mathcal{M}_{wm}$. It implies that $\mathcal{A}$’s success hinges on the possibility that embedding her watermark could overwrite (i.e., remove) $\mathcal{O}$’s watermark. In our series of experiments (cf. Appendix D.1), we observe that $\mathcal{O}$’s watermark cannot be overwritten after embedding $\mathcal{A}$’s watermark, which shows Genie’s robustness against ownership piracy attacks.
5.6 Dealing with adaptive attackers
Our results demonstrate Genie’s robustness against classical watermark removal techniques, including model extraction (cf. §5.4.15.4.2), model overwriting (cf. §5.4.35.4.5), and piracy attack (cf. §5.5). While these tests provide valuable insights into Genie’s overall robustness, it is crucial to evaluate its performance against newer attacks; in particular an adaptive attack, where $\mathcal{A}$ would design and implement an attack specifically tailored to Genie. Consequently, we evaluate Genie under stronger assumptions than previously considered attacks (e.g., in model extraction, access to resources such as $\mathcal{G}$ and $\mathcal{M}_{wm}$ was restricted). In the adaptive attack setting that we consider, access to both $\mathcal{G}$ and $\mathcal{M}_{wm}$ will be given to $\mathcal{A}$, and robustness of Genie’s watermark will be evaluated under these harsher assumptions.
To simulate real-world scenario, we assume that $\mathcal{A}$ accesses $\mathcal{M}_{own}$ (i.e., $\mathcal{M}_{wm}$) through an MLaaS system, querying $\mathcal{M}_{own}$ using only $\mathcal{V}$. It is analogous to the standard assumption of the user of an MLaaS being oblivious to its underlying complexities (e.g., a user avails the service of a recommendation MLaaS system using only node IDs, i.e., $\mathcal{V}$, while being oblivious to the underlying $(\mathbf{A},~{}\mathbf{X})$). We summarise the state of an adaptive $\mathcal{A}$ as follows: (1) $\mathcal{A}$ has access to $\mathcal{M}_{adv}$ (i.e., $\mathcal{M}_{wm}$) apart from $\mathcal{O}$’s MLaaS; (2) $\mathcal{A}$ understands Genie and knows that $\mathcal{M}_{adv}$ has been watermarked using Genie; (3) $\mathcal{A}$ knows the watermarking rate used (viz., $\alpha_{nr}$ or $\alpha_{sg}$); and (4) $\mathcal{A}$ knows the original graph $\mathcal{G}$.
The only information which $\mathcal{A}$ cannot infer from knowing Genie are $\mathcal{D}_{wm}$ and $\mathcal{G}_{wm}$, which are secret, since they are created by random sampling. We discuss the viability of an adaptive attack that exploits this information for node representation and subgraph-based methods of Genie as follows:
Node representation-based methods: In hopes to break Genie (i.e., to remove the watermark from $\mathcal{M}_{adv}$), $\mathcal{A}$ may attempt to guess links present in $\mathcal{D}_{wm}$ and construct $\mathcal{G}_{wm}$ by continuously querying $\mathcal{O}$’s MLaaS system. If successful, $\mathcal{A}$ can compare $\mathcal{G}_{wm}=(\mathcal{V},~{}\mathcal{E}_{wm})$ with $\mathcal{G}=(\mathcal{V},~{}\mathcal{E})$ to get the randomly sampled watermark links $\mathcal{S}_{wm}=(\mathcal{E}\backslash\mathcal{E}_{wm})\cup(\mathcal{E}_{wm}\backslash\mathcal{E})$ and then fine-tune $\mathcal{M}_{adv}$ with the labels opposite to $\mathcal{S}_{wm}$, potentially removing the watermark.
To defend Genie against such an attack, $\mathcal{O}$ can design the MLaaS system to invert the output whenever a user attempts to query the links in $\mathcal{S}_{wm}$. Consequently, any attempt to reconstruct $\mathcal{G}_{wm}$ by querying $\mathcal{O}$’s MLaaS system would only result in reconstruction of $\mathcal{G}$ instead of $\mathcal{G}_{wm}$. To conclude, such a defense closes all doors for $\mathcal{A}$ to guess $\mathcal{G}_{wm}$, thereby protecting $\mathcal{O}$ against such an adaptive $\mathcal{A}$.
Subgraph-based methods: If a link present in $\mathcal{D}_{wm}$ is queried to $\mathcal{O}$’s MLaaS system, the returned output will not be watermarked, i.e., the MLaaS system will classify the link correctly. It is because during inference: (1) $\mathcal{M}_{own}$ constructs the $k$-hop subgraph $\mathcal{G}_{k}$ surrounding the link; and (2) performs binary classification of $\mathcal{G}_{k}$. Since $\mathcal{G}_{k}$’s node feature vectors $\mathbf{x}_{v}$ have not been replaced with the secret watermark vector $\mathbf{w}$ present in $\mathcal{D}_{wm}$, $\mathcal{M}_{own}$ will output the correct prediction of the link. Therefore, the guessing of links present in $\mathcal{D}_{wm}$ by querying $\mathcal{O}$’s MLaaS system is infeasible. Consequently, the exploitation of $\mathcal{D}_{wm}$’s knowledge is not possible, in case of subgraph-based methods.
@@ 5.7 Efficiency
The computational efficiency of a watermarking scheme is crucial, as it directly impacts its cost-effectiveness and practical adoption. In Genie, the computational overhead is primarily determined by the size of the trigger set, as each epoch requires separate backpropagation for both the training set and the trigger set. Table A.2 illustrates the computational overhead in terms of $\alpha_{sg}$ and $\alpha_{nr}$, representing the relative sizes of the trigger sets for subgraph-based and node representation-based methods, respectively. Our analysis reveals that the maximum $\alpha_{sg}$ is 40% and the maximum $\alpha_{nr}$ is 15%, both occurring in smaller datasets. Notably, for large datasets, both $\alpha_{sg}$ and $\alpha_{nr}$ remain below 4%, making Genie scalable with dataset size. It demonstrates Genie’s computational overhead is reasonable, especially for large-scale real-world applications.
@@ 6 Conclusion
Despite the tremendous success of GNNs in learning graph-structured data, protecting trained GNN models from model-stealing attacks is a critical issue. Existing GNN watermarking schemes focus either on node or graph classification tasks. In this paper, we propose Genie, a backdoor-based watermarking scheme for GNNs tailored to LP task. We design Genie for node representation-based and subgraph-based methods of LP. Our exhaustive evaluations show that Genie satisfies all the watermarking requirements (cf. §3). At the same time, Genie is robust against SotA watermark removal techniques, model extraction attacks and an adaptive attack. The statistical assurance given by the OD procedure confirm a close to zero probability of misclassification in Genie. Given the salient features of Genie, especially its efficiency in accommodating large datasets and performance on diverse datasets of varying scales and sparsities, we claim that Genie is suitable for real-world application. We hope our work sets new research directions and benchmarks in the domain of model watermarking.
Appendix A Experiment setup details
A.1 Dataset description
USAir [45] is a network of US Airlines. NS [46] is a collaboration network of researchers in network science. Yeast [41] is a protein-protein interaction network in yeast. C.ele [44] is a neural network of C.elegans. Power [44] is an electrical grid network of the western US. arXiv [47] is a collaboration network of arXiv Astro Physics from the popular Stanford SNAP dataset library. PPI [48] is a protein-protein interaction network from BioGRID database. The dataset statistics are given in Table A.1.
TABLE A.1: Dataset statistics
Dataset Nodes Edges
USAir 332 2,126
NS 1,589 2,742
Yeast 2,375 11,693
C.ele 297 2,148
Power 4,941 6,594
arXiv 18,772 198,110
PPI 3,890 76,584
A.2 Model setup
SEAL: We use DGCNN as the GNN engine of SEAL. We use the default setting of DGCNN, i.e., four convolutional layers (32, 32, 32, 1 channels), a SortPooling layer (with $k=0.6$), two 1-D convolution layers (with 16, 32 output channels), and a 128-neuron dense layer. We train our models for a total of 50 epochs (for both training with or without a watermark). We use a learning rate of 0.0001.
GCN, SAGE: We use a 3-layer GCN and GraphSAGE model with a hidden layer of dimension 256. We use a 3-layer MLP for downstream binary classification with 256 hidden layer neurons. We train our models for a total of 400 epochs (for both training with or without a watermark). We use a learning rate of 0.001.
Neo-GNN: We use a 3-layer GCN with a hidden channel dimension of 256 as the GNN engine of Neo-GNN. We use a 3-layer MLP for downstream binary classification with 256 hidden layer neurons. We train our models for a total of 400 epochs (for both training with or without a watermark). We use a learning rate of 0.001.
Both $\mathcal{L}_{train}$ and $\mathcal{L}_{wm}$ use the same loss function (i.e., negative log likelihood) and optimizer (i.e., Adam) with the same learning rate. TABLE A.2 lists our watermarking rate for each dataset and respective model.
TABLE A.2: Watermarking rate (i.e., $\alpha_{sg}$ for SEAL and $\alpha_{nr}$ for GCN, GraphSAGE, and NeoGNN) used in our experiments.
@@ A.2 Model setup
SEAL 30 30 35 20 40 3 4
NeoGNN 10 15 10 4 5 2 4
Appendix B DWT: Theoretical analysis & proof
DWT uses KDE to calculate thresholds. The idea is to estimate the distribution of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ from as few data samples as possible. Once the distributions have been estimated, a threshold can be chosen that does not intersect with either of the distributions (i.e., a threshold that minimizes FPR and FNR). However, using fewer samples in KDE may lead to unfaithful estimation of distribution, one that may not be representative of the true distribution. Therefore, to faithfully estimate and represent the one-dimensional distributions, we use the result from [42], which states the following: at least 4 samples are required to bound the relative Mean Squared Error (MSE) of a one-dimensional KDE to 0.1. Formally, DWT can be broken down into a three-step process: Estimation, Sampling, and Thresholding.
  1. 1.
    Estimation: This step involves estimating distributions of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ using KDE. Let the estimated distribution of $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{clean}}$ and $AUC_{\mathcal{D}_{wm}}^{\mathcal{M}_{wm}}$ be $\mathcal{P}_{clean}$ and $\mathcal{P}_{wm}$, respectively. As mentioned before, at least 4 samples from both distribution are required to bound the relative MSE to 0.1. The bandwidth is set generously using Silverman’s rule of thumb that upper/lower bounds $\mathcal{P}_{clean}$ and $\mathcal{P}_{wm}$ respectively to reduce random error by introducing a systemic bias into the estimates [42].
  2. 2.
    Sampling: This step involves drawing multiple ($m\geq 3$) random sample of size $n$ [43] from both $\mathcal{P}_{clean}$ and $\mathcal{P}_{wm}$. We call $n$ as the sampling rate parameter of DWT.
  3. 3.
    Thresholding: This step involves selecting a threshold $t$ that minimizes FPR and FNR. Given $\mathcal{P}_{clean}$ and $\mathcal{P}_{wm}$ are sufficiently separated and non-overlapping [^7] The non-trivial ownership property along with its extremely low $p$-value (demonstrated in §4.3.1) show $\mathcal{P}_{clean}$ and $\mathcal{P}_{wm}$ will always be well-separated/non-overlapping. See Appendix C for more $p$-values., there would exist a (range of) $t$ for which the observed FPR and FNR is 00. Then, we select a $t$ such that the observed FPR and FNR is 00 across $m$ draws of the random sample.
To verify the correctness of DWT, we scrutinize each of the above steps as follows:
$\bullet$ The correctness of Estimation step is established by the 0.1 bound of relative MSE given in [42] with respect to the number of samples. To get a general bound on MSE with respect to the number of samples, we note from works [65, 66] that in KDE, if $f(x)$ is to be estimated and $\hat{f}(x)$ is the estimated KDE, then the optimal MSE at point $x$ has an approximate scaling law in the following form:
$\text{MSE}(\hat{f}(x))\sim C_{d}\cdot n^{-\frac{4}{4+d}},$ (B.1)
where $d$ is the dimension, $n$ is sample size, and $C_{d}$ is the dimension dependent constant. Therefore, if a sample size $n_{0}$ is known to achieve a certain MSE level $\epsilon_{0}$, then using Eq. B.1, one can derive the estimate of the sample size $n_{1}$ to achieve a different MSE $\epsilon_{1}$ by Eq. B.2 as follows:
$n_{1}\approx n_{0}\times\left(\frac{\epsilon_{0}}{\epsilon_{1}}\right)^{\frac{4+d}{4}}$ (B.2)
$\bullet$ Since Sampling step involves only sampling from probability distributions $\mathcal{P}_{clean}$ and $\mathcal{P}_{wm}$, there are no correctness considerations in it. However, there are practical consideration about the complexity of the sampling procedure in achieving desired FPR and FNR. Theoretically, the system can achieve any desired $\frac{1}{n}$ probability bound by varying the sampling parameter $n$. However, since $n$ is the number of samples drawn, choosing a large $n$ could be computationally expensive.
$\bullet$ Finally, as a part of correctness of Thresholding step, we give a proof of the result stated in §4.3.2. Before giving the proof, we note that each draw of the random sample of size $n$ follows a Binomial Distribution, $\mathrm{Binomial}(n,p)$, where $p$ is the probability of misclassification (viz., FPR, FNR). Here, each of the $m$ draws of the random sample can be considered as a block, where 0 misclassification has happened (i.e., 0 observed FPR and FNR). In what follows, we present the result as a theorem and its proof.
Theorem. Let $X_{j}\sim\mathrm{Binomial}(n,p)$ be the number of “misclassifications” observed in the $j$-th block of $n$ Bernoulli trials, for $j=1,\dots,m$. If $X_{j}=0$ for all $j$ and
$m\;\geq\;\bigl{\lceil}-\ln(1-\gamma)\bigr{\rceil},$ (B.3)
then with confidence at least $\gamma$, we have $p<\frac{1}{n}$.
Proof. We first note that
$\Pr(X_{j}=0)\;=\;(1-p)^{\,n}.$ (B.4)
Hence, observing $X_{j}=0$ for all $j=1,\dots,m$ yields
$\Pr(X_{1}=0,\ldots,X_{m}=0)\;=\;\bigl{(}(1-p)^{n}\bigr{)}^{m}\;=\;(1-p)^{\,mn}.$ (B.5)
If $p\geq\frac{1}{n}$, then
$(1-p)\;\leq\;e^{-p}\;\leq\;e^{-\tfrac{1}{n}},$ (B.6)
implying
$(1-p)^{\,mn}\;\leq\;\exp\bigl{(}-mnp\bigr{)}\;\leq\;e^{-m}.$ (B.7)
Thus,
Pr(all $\Pr~{}\!\bigl{(}\text{all zeros}\,\bigm{|}\,p\geq 1/n\bigr{)}\;\leq\;e^{-m}.$ (B.8)
If we choose $m$ such that
[…]
then the probability of observing zero events in all $m$ blocks despite $p\geq 1/n$ is at most $1-\gamma$. Equivalently, whenever we do observe zero events in all $m$ blocks, it follows with confidence at least $\gamma$ that $p<1/n$.  $\Box$
Therefore, to achieve $95\%$ confidence interval, i.e., $\gamma=0.95$, we require only $m=\lceil-\ln(1-0.95)\rceil=\lceil 2.997\rceil=3$ draws, which is often stated as the rule-of-three [67]. Similarly, 5 draws will lead to $\gamma=1-e^{-5}=0.9973$, i.e., $99.73\%$ confidence interval of $p\in[0,\frac{1}{n})$.
Appendix C Statistical assurance of non-trivial ownership in Genie